# Analysis of cohort stepped wedge cluster-randomized trials with non-ignorable dropout via joint modeling

### v1

@@ Abstract
Stepped wedge cluster-randomized trial (CRTs) designs randomize clusters of individuals to intervention sequences, ensuring that every cluster eventually transitions from a control period to receive the intervention under study by the end of the study period. The analysis of stepped wedge CRTs is usually more complex than parallel-arm CRTs due to potential secular trends that result in changing intra-cluster and period-cluster correlations over time. A further challenge in the analysis of closed-cohort stepped wedge CRTs, which follow groups of individuals enrolled in each period longitudinally, is the occurrence of dropout. This is particularly problematic in studies of individuals at high risk for mortality, which causes non-ignorable missing outcomes. If not appropriately addressed, missing outcomes from death will erode statistical power, at best, and bias treatment effect estimates, at worst. Joint longitudinal-survival models can accommodate informative dropout and missingness patterns in longitudinal studies. Specifically, within the joint longitudinal-survival modeling framework, one directly models the dropout process via a time-to-event submodel together with the longitudinal outcome of interest. The two submodels are then linked using a variety of possible association structures. This work extends linear mixed-effects models by jointly modeling the dropout process to accommodate informative missing outcome data in closed-cohort stepped wedge CRTs. We focus on constant intervention and general time-on-treatment effect parametrizations for the longitudinal submodel and study the performance of the proposed methodology using Monte Carlo simulation under several data-generating scenarios. We illustrate the joint modeling methodology in practice by reanalyzing data from the ‘Frail Older Adults: Care in Transition’ (ACT) trial, a stepped wedge CRT of a multifaceted geriatric care model versus usual care in 35 primary care practices in the Netherlands.
@@
closed-cohort designs, stepped wedge designs, joint longitudinal-survival modeling, Monte Carlo simulation, informative attrition
@@ 1 Introduction
Cluster-randomized trials (CRTs) are popular study designs used to evaluate interventions delivered to a group of individuals, such as clinics, hospitals, or villages. CRTs are adopted when the cluster is the intervention target, when individual randomization is infeasible, or when the risk for contamination is high 1. In the most straightforward design, clusters in CRTs are randomized in parallel, with each cluster randomly assigned to either intervention or control with no crossover. An alternative CRT design uses the so-called stepped wedge randomization, with clusters randomized to crossover from a control condition at different times, i.e., randomizing when the intervention condition will start 2, 3. In this design, each cluster transitions to the intervention at a different point in time, and by the end of the trial, all clusters have received the intervention under study. Benefits and tradeoffs of stepped wedge designs are discussed in detail elsewhere 4, 3; among others, they overcome difficulties in rolling out the intervention to all study participants at the same time, they allow every study participant to be exposed to the study intervention(s), and they can increase statistical power as each cluster serves as its own control. There are several types of stepped wedge designs, including open, cross-sectional, and closed-cohort, each referring to how participants are followed in the study 5, 6. The work focuses on inferential methods for non-ignorable missing outcome data due to death in closed-cohort designs, where the outcomes of individuals are measured repeatedly over the study period (Figure 1).
@@
The loss of study participants in longitudinal studies is a major challenge that, if not adequately handled, can introduce bias in the analysis and affect the identification of true treatment effects. This is especially relevant when the loss to follow-up is correlated to socioeconomic position or poor health, which can correlate with a higher risk of dropping out 7. For instance, a recent review of nursing home CRTs by Poupin et al. found that a median of 19.5% study participants did not complete the entire study protocol, mostly because of death 8.
@@
Dropout in longitudinal studies is often ignored or inadequately examined to assess the potential impact of missing outcome data (and approaches to handling missing data) on study conclusions 8. In this manuscript, we approach this problem by explicitly modeling the dropout process together with the longitudinal outcome within the shared random-effects modeling framework. Other approaches, such as locating and re-engaging study participants who were lost to follow-up, have been proposed in the literature 9; this is, however, not always possible, particularly when dropout is due to the death of the study participant. Our approach is tailored to the settings of closed-cohort stepped wedge CRTs but could be extended to parallel and cluster-crossover CRTs (and other multilevel settings), when repeated measurements are collected over multiple time periods, as well.
@@
It is known in the literature that dropout from the study, per se, does not necessarily introduce bias in the analysis. For instance, if dropout is missing completely at random (MCAR), then the dropout process is ignorable and only leads to loss of efficiency. In studies where MCAR is expected, inflation of the planned sample size is a valuable strategy to prevent loss of power, and standard methods for the analysis of longitudinal data can be used without bias 10, 8, 11, 12. Mixed-effects models estimated on the available data are recommended in these settings, as their maximum likelihood estimators are robust to missing at random (MAR); conversely, generalized estimating equations (GEE) estimators are only unbiased when data is MCAR, but a weighted GEE can be used to overcome this limitation 11, 13, 12. When dropout is not MCAR nor MAR, then it is said to be missing not at random (MNAR), informative, or non-ignorable: in these settings, the selection bias over time induced by differential attrition rates (e.g., between treatment arms) can lead to bias and significant additional heterogeneity, as illustrated with simulated data in Figure 2. For example, in the MNAR settings, attrition rates may depend on unobserved characteristics of the longitudinal outcome, such as the current true value or the slope of change. Note that dropout from the study may be informative even with equal attrition rates between treatment arms, i.e., if dropout occurs for different reasons, such as toxicity in the treatment arm and lack of perceived benefit in the control arm. This leads to a form of informative censoring that is cryptic and hard to identify 14. The difference between MCAR, MAR, and MNAR settings in randomized trials is further discussed, among other methodological considerations, by Thakur et al. 15
@@
Figure 2: Illustrative example of biased longitudinal trajectories in the settings of informative dropout, using simulated data. The example assumes that study participants drop out when the longitudinal outcome exceeds the value of five, e.g., representing a hypothetical scenario where participants have improved enough that they no longer perceive the benefits of continuing the study.
The issue of informative dropout in trials with longitudinal outcomes, and the role of joint modeling specifically, has been previously studied in the literature. Rouanet et al. summarized different approaches for modeling longitudinal outcomes and highlighted how joint longitudinal-survival models can accommodate MNAR settings, while mixed models and (potentially weighted) GEE models cannot 12. Kolamunnage-Dona et al. illustrated the joint modeling approach in randomized trials (without any clustering) with competing causes of dropout, highlighting the value of this approach for evaluating the sensitivity of conclusions to assumptions regarding missing data mechanisms 16. Cuer et al. compared linear mixed models and joint longitudinal-survival models to show their differences in terms of outputs, interpretation, and underlying modeling assumptions 17. In further work, they compared standard joint longitudinal-survival models with competing risks joint models (e.g., those discussed by Kolamunnage-Dona et al. as well 16) both in practice and via simulation, showing that the standard model and the competing risks model performed equally well when the risk of dropout was the same whatever the cause 18. Moreover, they showed that the standard joint model led to biased results when the risk of dropout differed between causes; this was not the case for the competing risks joint model 18. Hogan et al. developed a mixture-modeling approach for longitudinal data with outcome-dependent dropout where the functional dependence between covariate effects and dropout time can be left unspecified 19; their approach can be considered to be an extension of selection models 20, 21. Wang and Chinchilli developed a shared parameters model for the settings of crossover designs with informative dropout 22. Specifically, they combine mixed-effects models with a model for dropout based on fully parametric discrete hazard models, and they implement maximum likelihood estimation to estimate the model parameters. Moreover, they extended their approach to binary longitudinal outcomes and showed that GEE models and generalized linear mixed models with pseudo-likelihood estimation for complete and incomplete data were robust to informative dropout, highlighting that, in their simulation settings, the impact from the missing data on the model estimates was mild with a binary outcome 23. Finally, Poupin et al. suggested that open-cohort designs could be used to reduce the impact of potentially informative attrition, as long as (1) the intervention and the main cause of attrition are independent and (2) the intervention has a cluster-specific effect (if any) 24. This is, therefore, not feasible with interventions that are expected to have an individual effect, such as medications or treatments administered at the subject level, and is beyond the scope of this manuscript. Despite this existing work, there has been limited research on methods for addressing informative attrition in closed-cohort stepped wedge CRTs, where the exogenous, time-dependent treatment indicator can affect both the attrition rate and the longitudinal outcome. In addition, recent studies have indicated the importance of accounting for time-dependent treatment effects in the longitudinal modeling in stepped wedge CRTs, and it remains of interest to develop methods in the presence of both time-dependent treatment effects and informative attrition 25, 26, 27.
@@
The rest of this manuscript is organized as follows. In Section 2, we introduce the mixed modeling parametrization for closed-cohort stepped wedge CRTs that we focus on throughout this work. In Section 3, we extend the mixed modeling approach to the joint longitudinal-survival framework, including model formulation, likelihood, and estimation. In Section 4, we compare the two approaches using Monte Carlo simulation methods. In Section 5, we re-analyze data from the Frail Older Adults: Care in Transition (ACT) trial, a pragmatic stepped wedge CRT of a multifaceted geriatric care model versus usual care in 35 primary care practices in the Netherlands, using both the traditional mixed modeling and the joint modeling approach. In Section 6, we discuss extensions to the proposed framework. Finally, we conclude the manuscript with a discussion in Section 7.
@@ 2 Mixed-effects models: a recap
Closed-cohort stepped wedge trials are routinely analyzed using mixed-effects models 3. For this project, we focus on the following mixed-effects model:
@@
where $Y_{ik}(t)$ is a continuous outcome measured at time $t$ for the ith cluster during the jth discrete period and on the kth study participant. $X_{i}(t)$ is a binary covariate denoting whether the ith cluster received treatment during the discrete calendar period $(T_{j-1},T_{j}]$, $\delta$ is the treatment effect, and $\beta_{j}$ are period-effect coefficients. We assume that $\varepsilon\sim N(0,\sigma^{2}_{\varepsilon})$ is residual (measurement) error, and we include two independent random effects, $\alpha\sim N(0,\sigma^{2}_{\alpha})$ and $\phi\sim N(0,\sigma^{2}_{\phi})$, representing a cluster-, and a subject-specific random intercept, respectively. This model is sometimes referred to as a nested exchangeable model, with a constant intervention effect; it represents an extension of the Hussey and Hughes model introduced by Baio et al. 28. A nested exchangeable model implies the following correlation structure:
@@
assuming the intervention was introduced during period $s$. This is illustrated in Figure 3. This model estimates a potentially distinct treatment effect per level of the exposure time (duration of the treatment and is different from the calendar time), and addresses time-dependent treatment effect heterogeneity, which may arise due to the learning, weakening, or delayed effect due to implementation of the intervention 3, 25, 26, 27.
Figure 3: Example of a closed-cohort stepped wedge trial with $i=4$ clusters and $j=5$ time periods, with a general time on treatment effect $\delta_{j-s}$; grey boxes represent intervention periods, while white boxes represent control periods.

### v3

@@ Abstract
Stepped wedge cluster-randomized trial (CRTs) designs randomize clusters of individuals to intervention sequences, ensuring that every cluster eventually transitions from a control period to receive the intervention under study by the end of the study period. The analysis of stepped wedge CRTs is usually more complex than parallel-arm CRTs due to more complex intra-cluster correlation structures. A further challenge in the analysis of closed-cohort stepped wedge CRTs, which follow groups of individuals enrolled in each period longitudinally, is the occurrence of dropout. This is particularly problematic in studies of individuals at high risk for mortality, which causes non-ignorable missing outcomes. If not appropriately addressed, missing outcomes from death will erode statistical power, at best, and bias treatment effect estimates, at worst. Joint longitudinal-survival models can accommodate informative dropout and missingness patterns in longitudinal studies. Specifically, within the joint longitudinal-survival modeling framework, one directly models the dropout process via a time-to-event submodel together with the longitudinal outcome of interest. The two submodels are then linked using a variety of possible association structures. This work extends linear mixed-effects models by jointly modeling the dropout process to accommodate informative missing outcome data in closed-cohort stepped wedge CRTs. We focus on constant intervention and general time-on-treatment effect parametrizations for the longitudinal submodel and study the performance of the proposed methodology using Monte Carlo simulation under several data-generating scenarios. We illustrate the joint modeling methodology in practice by reanalyzing data from the ‘Frail Older Adults: Care in Transition’ (ACT) trial, a stepped wedge CRT of a multifaceted geriatric care model versus usual care in 35 primary care practices in the Netherlands.
@@
closed-cohort designs, stepped wedge designs, joint longitudinal-survival modeling, Monte Carlo simulation, informative attrition, time-varying treatment effect
@@
The published version of this manuscript can be accessed online at the following URL: http://dx.doi.org/10.1002/sim.10347
@@ 1 Introduction
Cluster-randomized trials (CRTs) are popular study designs used to evaluate interventions delivered to a group of individuals, such as clinics, hospitals, or villages. CRTs are adopted when the cluster is the intervention target, when individual randomization is infeasible, or when the risk for contamination is high. 1 In the most straightforward design, clusters in CRTs are randomized in parallel, with each cluster randomly assigned to either intervention or control with no crossover. An alternative CRT design uses the so-called stepped wedge randomization, with clusters randomized to crossover from a control condition at different times, i.e., randomizing when the intervention condition will start. 2, 3 In this design, each cluster transitions to the intervention at a different point in time, and by the end of the trial, all clusters have received the intervention under study. Benefits and tradeoffs of stepped wedge designs are discussed in detail elsewhere 4, 5, 3; among others, they overcome difficulties in rolling out the intervention to all study participants at the same time, they allow every study participant to be exposed to the study intervention(s), and they can, in some settings, increase statistical power as each cluster serves as its own control. For instance, Hemming and Taljaard compared the relative efficiency of parallel CRTs and stepped wedge CRTs and concluded that statistical power depends on within-cluster correlation and cluster size, with stepped wedge CRTs being more efficient than parallel CRTs when the intra-cluster correlation is high. 5 Moreover, there are several types of stepped wedge designs, including open, cross-sectional, and closed-cohort, each referring to how participants are followed in the study. 6, 7 The work focuses on inferential methods for non-ignorable missing outcome data due to death in closed-cohort designs, where the outcomes of individuals are measured repeatedly over the study period (Figure 1).
@@
The loss of study participants in longitudinal studies is a major challenge that, if not adequately handled, can introduce bias in the analysis and affect the identification of true treatment effects. This is especially relevant when the loss to follow-up is correlated to socioeconomic position or poor health, which can correlate with a higher risk of dropping out. 8 For instance, a recent review of nursing home CRTs by Poupin et al. found that a median of 19.5% study participants did not complete the entire study protocol, mostly because of death. 9
@@
Dropout in longitudinal studies is often ignored or inadequately examined to assess the potential impact of missing outcome data (and approaches to handling missing data) on study conclusions. 9 In this manuscript, we approach this problem by explicitly modeling the dropout process together with the longitudinal outcome within the shared random-effects modeling framework. Other approaches, such as locating and re-engaging study participants who were lost to follow-up, have been proposed in the literature 10; this is, however, not always possible, particularly when dropout is due to the death of the study participant. Our approach is tailored to the settings of closed-cohort stepped wedge CRTs but could be extended to parallel and cluster-crossover CRTs (and other multilevel settings), when repeated measurements are collected over multiple time periods, as well.
@@
It is known in the literature that dropout from the study, per se, does not necessarily introduce bias in the analysis. For instance, if dropout corresponds to outcome missing completely at random (MCAR), then the dropout process is ignorable and only leads to loss of efficiency. In studies where MCAR is expected, inflation of the planned sample size is a convenient strategy to restore study power, and standard methods for the analysis of longitudinal data can be used without bias. 11, 9, 12, 13 Mixed-effects models estimated on the available data are recommended in these settings, as their maximum likelihood estimators are robust to missing at random (MAR); conversely, generalized estimating equations (GEE) estimators are only unbiased when data is MCAR, but a weighted GEE can be used to overcome this limitation under MAR. 12, 14, 13 When dropout is not MCAR nor MAR, then it is said to be missing not at random (MNAR), informative, or non-ignorable: in these settings, the selection bias over time induced by differential attrition rates (e.g., between treatment arms) can lead to bias. We illustrate this with a simulated example in Figure 2, where we randomized 50 patients to a binary treatment (or control) and assume a non-null treatment effect (i.e., treated subjects improve/increase their longitudinal outcome over time). We also assume that study subjects leave the study once their outcome exceeds the value of five, e.g., because they have improved enough to no longer perceive the benefits of continuing the study. In this setting, we would underestimate the benefit of the treatment if we were ignoring this informative dropout (red dotted line) compared to the true underlying longitudinal trajectory (red solid line), with an estimated linear slope in the observed treatment arm of 0.64 (95% C.I.: 0.44, 0.84) instead of a true slope of 1. Note that, in the MNAR settings, attrition rates may depend on unobserved characteristics of the longitudinal outcome, such as the current true value or the slope of change, and that dropout from the study may be informative even with equal attrition rates between treatment arms, i.e., if dropout occurs for different reasons, such as toxicity in the treatment arm and lack of perceived benefit in the control arm. This leads to a form of informative censoring that is cryptic and hard to identify. 15 The difference between MCAR, MAR, and MNAR settings in randomized trials is further discussed, among other methodological considerations, by Thakur et al. 16
@@
Figure 2: Illustrative example of biased longitudinal trajectories in the settings of informative dropout, using simulated data. The example assumes that study participants drop out when the longitudinal outcome exceeds the value of five, e.g., representing a hypothetical scenario where participants have improved enough to no longer perceive the benefits of continuing the study.
@@
The rest of this manuscript is organized as follows. In Section 2, we review and discuss the current literature on informative dropout in trials with a longitudinal outcome. In Section 3, we introduce the mixed modeling parametrization for closed-cohort stepped wedge CRTs that we focus on throughout this work. In Section 4, we extend the mixed modeling approach to the joint longitudinal-survival framework, including model formulation, likelihood, and estimation. In Section 5, we compare the two approaches using Monte Carlo simulation methods. In Section 6, we re-analyze data from the Frail Older Adults: Care in Transition (ACT) trial, a pragmatic stepped wedge CRT of a multifaceted geriatric care model versus usual care in 35 primary care practices in the Netherlands, using both the traditional mixed modeling and the joint modeling approach. In Section 7, we discuss extensions to the proposed framework. We conclude the manuscript with a discussion in Section 8.
2 Literature review and objectives
The issue of informative dropout in trials with longitudinal outcomes, and the role of joint modeling specifically, has been previously studied in the literature. Rouanet et al. summarized different approaches for modeling longitudinal outcomes and highlighted how joint longitudinal-survival models can accommodate MNAR settings, while mixed models and (potentially weighted) GEE models cannot. 13 Kolamunnage-Dona et al. illustrated the joint modeling approach in randomized trials (without any clustering) with competing causes of dropout, highlighting the value of this approach for evaluating the sensitivity of conclusions to assumptions regarding missing data mechanisms. 17 Cuer et al. compared linear mixed models and joint longitudinal-survival models to show their differences in terms of outputs, interpretation, and underlying modeling assumptions. 18 In further work, they compared standard joint longitudinal-survival models with competing risks joint models (e.g., those discussed by Kolamunnage-Dona et al. as well 17) both in practice and via simulation, showing that the standard model and the competing risks model performed equally well when the risk of dropout was the same whatever the cause. 19 Moreover, they showed that the standard joint model led to biased results when the risk of dropout differed between causes; this was not the case for the competing risks joint model. 19 Hogan et al. developed a mixture-modeling approach for longitudinal data with outcome-dependent dropout where the functional dependence between covariate effects and dropout time can be left unspecified 20; their approach can be considered to be an extension of selection models. 21, 22 Wang and Chinchilli developed a shared parameters model for the settings of crossover designs with informative dropout. 23 Specifically, they combine mixed-effects models with a model for dropout based on fully parametric discrete hazard models, and they implement maximum likelihood estimation to estimate the model parameters. Moreover, they extended their approach to binary longitudinal outcomes and showed that GEE models and generalized linear mixed models with pseudo-likelihood estimation for complete and incomplete data were robust to informative dropout, highlighting that, in their simulation settings, the impact from the missing data on the model estimates was mild with a binary outcome. 24
We now describe and compare in more detail the three main approaches that have been proposed for modeling MNAR outcomes: pattern mixture models, selection models, and joint longitudinal-survival models. First, pattern mixture models factorize the joint distribution of the longitudinal outcome and missingness mechanism into the probability of missing data and the conditional distribution of the outcome given missingness. One such example for analyzing continuous outcomes in longitudinal cluster randomized trials was developed by Fiero et al., where they combined pattern mixture models with multilevel multiple imputation to create MNAR imputed values. 25 This approach provides a viable strategy for performing sensitivity analyses under certain user-defined assumptions regarding the missing data mechanism. Demirtas and Schafer discuss random coefficient pattern mixture models, that is, random-effects models for a certain longitudinal outcome that include summaries of the dropout mechanism as fixed effects. 26 In their work, they illustrate how alternative random coefficient pattern mixture models that fit the data equally well may lead to very different estimates for the parameters of interest, highlighting how even minor model misspecification can introduce large biases. Second, selection models factorize the joint distribution of longitudinal outcome and missingness the other way around, as a model for the study outcome (the substantive analysis) times the conditional probability of missingness given the outcome; specifically, the Diggle-Kenward selection model combines a growth curve model (as the substantive model) with a set of logistic or probit regression models (as the conditional distribution) that directly relate the missing data probability to the outcome variable at current and/or previous time points. 21 An extension of selection models is given by Hogan et al. 20 Moreover, some of these approaches have been previously compared via Monte Carlo simulation, concluding that selection models outperformed pattern mixture models — especially as sample size and dropout rate increased. 27 Third, the shared random effects, joint modeling approach introduces a common set of latent random effects that link the longitudinal outcome and the dropout mechanism. 28, 29, 30 The longitudinal and survival outcomes are then assumed to be conditionally independent (given the random effects) and are jointly estimated within a unified framework, with flexibility in terms of model formulations for the two submodels. An example is given by Wang and Chinchilli in the context of crossover trials. 23
The three approaches for MNAR outcomes described above have advantages and disadvantages. Pattern mixture models and selection models can have relatively simple formulations and are usually easier to estimate compared to joint longitudinal-survival models. However, these approaches require distributional assumptions for model parameters to be identifiable, and it has been shown that fitted parameters may vary widely under different missing data assumptions. 26, 25, 27 On the other end, shared random-effects joint models are computationally more challenging to estimate, and the population-level interpretation of the fitted coefficients for the outcome model is only valid in an immortal cohort. 12 Nonetheless, we choose to focus on this approach for a variety of reasons. First, joint longitudinal-survival models provide much more flexibility (in terms of possible model formulations) currently available in statistical software packages, with further improvements as software implementations develop. Second, this class of models is much closer to standard approaches for the analysis of stepped wedge CRTs based on mixed-effects models, 3 as they extend the usual methods by explicitly modeling the dropout process on top of the outcome model. Therefore, we believe these methods to be much more approachable by practitioners. Third, joint longitudinal-survival models do not significantly contaminate the conventional model-based target treatment effect parameters in stepped wedge CRTs, whereas other methods may do so without careful reparameterization.
Finally, ther design approaches to tackle the issue of informative dropout are also possible. For instance, Poupin et al. suggested that open-cohort designs could be used to reduce the impact of potentially informative attrition, as long as (1) the intervention and the main cause of attrition are independent and (2) the intervention has a cluster-specific effect (if any). 31 This is, therefore, not feasible with interventions that are expected to have an individual effect, such as medications or treatments administered at the subject level, and is beyond the scope of this manuscript. Despite this existing work, there has been limited research on methods for addressing informative attrition in closed-cohort stepped wedge CRTs, where the exogenous, time-dependent treatment indicator can affect both the attrition rate and the longitudinal outcome. In addition, recent studies have indicated the importance of accounting for time-dependent treatment effects in the longitudinal modeling in stepped wedge CRTs, and it remains of interest to develop methods in the presence of both time-dependent treatment effects and informative attrition. 32, 33, 34
@@ 3 Mixed-effects models: a recap
Closed-cohort stepped wedge trials are routinely analyzed using mixed-effects models. 3 For this project, we focus on the following linear mixed-effects model:
@@
where $Y_{ijk}(t)$ is a continuous outcome measured at time $t$ for the ith cluster during the jth discrete period and on the kth study participant. $X_{i}(t)$ is a binary covariate denoting whether the ith cluster received treatment during the discrete calendar period $(T_{j-1},T_{j}]$, $\delta$ is the treatment effect, and $\beta_{j}$ are period-effect coefficients. We assume that $\varepsilon\sim N(0,\sigma^{2}_{\varepsilon})$ is residual (measurement) error, and we include two independent random effects, $\alpha\sim N(0,\sigma^{2}_{\alpha})$ and $\phi\sim N(0,\sigma^{2}_{\phi})$, representing a cluster-, and a subject-specific random intercept, respectively. This model is sometimes referred to as a nested exchangeable random-effects model, with an immediate and constant intervention effect; it represents an extension of the Hussey and Hughes model introduced by Baio et al. 35 A nested exchangeable model implies the following correlation structure:
@@
assuming the intervention was introduced during period $s$. This is illustrated in Figure 3. This model estimates a potentially distinct treatment effect per level of the exposure time (duration of the treatment and is different from the calendar time), and addresses exposure-time-dependent treatment effect heterogeneity, which may arise due to the learning, weakening, or delayed effect due to implementation of the intervention. 3, 32, 33, 34
Figure 3: An example of a closed-cohort stepped wedge trial with $i=4$ clusters and $j=5$ time periods, with a general time on treatment effect $\delta_{j-s}$; grey boxes represent intervention periods, while white boxes represent control periods.